OpenStudy (anonymous):

Which expression is the first step you take when finding the quotient of 5/7 and 3/2?

OpenStudy (anonymous):

\[\frac{5}{7}\times \frac{2}{3}\]

OpenStudy (anonymous):

invert 3/2 then multiply